OpenStudy (scarlettthecamel):

what is the square root of 8

OpenStudy (anonymous):

2sqrt(2)

OpenStudy (anonymous):

\[\Large \sqrt{8}=\sqrt{4*2}=\sqrt{4} \times \sqrt{2}=2\sqrt{2}\]

OpenStudy (scarlettthecamel):

lost

OpenStudy (anonymous):

If you have a sqrt, then you can split it into its factors

OpenStudy (anonymous):

This site goes through how to simplify a radical really well http://www.regentsprep.org/regents/math/algebra/AO1/Lsimplify.htm

OpenStudy (confluxepic):

2.82

OpenStudy (anonymous):

Lol @confluxepic

OpenStudy (anonymous):

if you square root something, and it ends up being a decimals, you have to automatically know that you must simplify it.

OpenStudy (anonymous):

2.83*

OpenStudy (confluxepic):

lol

OpenStudy (scarlettthecamel):

2.83 so would it b just 2.8 or 3?

OpenStudy (scarlettthecamel):

pls guys

OpenStudy (scarlettthecamel):

don't mek me do this on me own

OpenStudy (scarlettthecamel):

cries

OpenStudy (scarlettthecamel):

i cri evritiem

OpenStudy (anonymous):

The answer is 2sqrt(2) if you want to be exact. If you want a rounded answer go with 2.828